OpenStudy (anonymous):

Logarithms: There exist a positive number k such that \[\log_{2}x+\log_{4}x+\log_{8}x=\log_{k}x\]for all positive real numbers x. If \[k=\sqrt[b]{a}\] where a,b are natural nos., the smallest possible value of (a+b) is equal to- Ans.75 (How????)

OpenStudy (anonymous):

@nitz

ganeshie8 (ganeshie8):

\(\huge log_{b^n}a = \frac{log_ba}{n}\)

ganeshie8 (ganeshie8):

equation simplifies to, \(\huge \frac{log_2x^{11}}{6} = log_kx\)

ganeshie8 (ganeshie8):

\(\huge log_2 x = \frac{6}{11}log_kx\)

ganeshie8 (ganeshie8):

\(\huge log_2 x = \frac{log_kx}{11/6}\) \(\huge log_2 x = log_{k^{(11/6)}}x\) \(\huge 2 = k^{11/6}\) \(\huge k = 2^{6/11}\) b = 11, a = 64

ganeshie8 (ganeshie8):

smallest possible value for (a, b) is (2^6, 11) = (64, 11) 6/11 = 12/22, so, next immediate value for (a, b) is (2^12, 11*2) = (4096, 22)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its done already